OpenStudy (anonymous):

xy = –25. What is the constant of variation for this relationship?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Is that direct, indirect, or joint variation?

OpenStudy (anonymous):

Errr, I dunno

OpenStudy (whpalmer4):

Okay, Direct variation: \(y = kx\) Indirect variation: \(y = k/x\) or \(xy = k\) Joint variation: \(z = kxy\)

OpenStudy (anonymous):

Indirect variation

OpenStudy (whpalmer4):

Very good. The constant of variation is?

OpenStudy (anonymous):

How do I find that

OpenStudy (whpalmer4):

Very easy. Write down "xy = -25" on a blank piece of paper. Look at the paper. Anything that isn't "x", "y", or "=" is the constant of variation :-)

OpenStudy (whpalmer4):

It's \(k\) in all of the formulas I gave you...

OpenStudy (anonymous):

So the constant of variation is -25?

OpenStudy (whpalmer4):

yes

OpenStudy (anonymous):

O.O Thats it?

OpenStudy (whpalmer4):

That's it! The key here was recognizing that \(xy=-25\) is the same as \(y = -25/x\)

OpenStudy (anonymous):

Thx